A man left a will stating that his pregnant wife had a daughter, he wants his daughter to have twicw as much as his wife. if his pregnant wife has a son, he wants his son to have half as much as his wife. After his death, his wife gave birth to twins; a boy and a girl. According to his will, what fraction of esate does each person get? Find geometric way to explain your thinking.
Answer by ewatrrr(24785)


Hi,
Let x represent the amount a son is left as part of the estate:
w =2x (twice the son)
D = 4x (twice the wife)
As there were twins.. estate is split 7 ways
Son = 1/7
Daughter = 4/7
Wife = 2/7
1/7 + 4/7 + 2/7 = the whole estate